OpenStudy (anonymous):

Simplify: Factor the expression. 6x^3y – 9x^2y^2 + 3x^2y

OpenStudy (skullpatrol):

Any ideas?

OpenStudy (skullpatrol):

$$6x^3y – 9x^2y^2 + 3x^2y$$

OpenStudy (anonymous):

-3xy+3x^2y= 0xy

OpenStudy (anonymous):

prime?

OpenStudy (skullpatrol):

What is that?

OpenStudy (anonymous):

oops

OpenStudy (anonymous):

hold on just a sec.

OpenStudy (anonymous):

the answer is 3x^2y

OpenStudy (skullpatrol):

Could you show me your work please?

OpenStudy (anonymous):

i do it in my head

OpenStudy (skullpatrol):

Orly?

OpenStudy (anonymous):

yes

OpenStudy (skullpatrol):

$$6x^3y – 9x^2y^2 + 3x^2y = 3x^2y$$ Is this your answer?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

its factoring

OpenStudy (skullpatrol):

And you did it ALL in your head?

OpenStudy (anonymous):

so \[3x ^{2}(2x-3y+1)\]

OpenStudy (anonymous):

YES

OpenStudy (anonymous):

no paper or pencils or calculators just my brain

OpenStudy (anonymous):

i swear

OpenStudy (skullpatrol):

$$6x^3y – 9x^2y^2 + 3x^2y$$ Start by factoring out the greatest monomial factor first.

OpenStudy (anonymous):

3x^2y is 9x^2y^2 factored

OpenStudy (skullpatrol):

$$6x^3y – 9x^2y^2 + 3x^2y= 3x^2y(2x-3y+1)$$

OpenStudy (anonymous):

yes i know, i figured it out in my head

OpenStudy (skullpatrol):

You missed a factor of y.

OpenStudy (anonymous):

no because only one of them has 2 y's if they all had y^2 then it would be 3x^2y^2 but its not because that doesn't go into 6x^3y or 3x^2y

OpenStudy (skullpatrol):

$$6x^3y – 9x^2y^2 + 3x^2y$$ All three terms have y in common.

OpenStudy (anonymous):

yes but only one y

OpenStudy (skullpatrol):

So factor it out.